85 is what percent 0f 60?
Answer by stanbon(75887)

85 is what percent 0f 60?
85/60 = x/100
17/12 = x/100
1700/12 =x
x= 141.6666667...
x=(141 and 2/3)%
Cheers,
Stan H.